### v1

# Winding quotients for virtual period maps of rank 1

Abstract.
We illustrate a rank 1 model of virtual period maps and their associated winding quotient, where the winding quotient is a new phenomenon appeared in a recent study of virtual period maps and it requires a reformulation of the classical inversion problem for the period maps. We answer to the new inversion problem by introducing the $q$-multiplicatively periodic function, whose pull-back to the winding covering space is the Weierstrass $\wp$-function up to a correction by Eisenstein series $E_{2}$, and which was used in recent mathematical physics.
2010 Mathematics Subject Classification:
30c99 (primary) 33E05 (secondary)
1. Introduction
The classical period map for a primitive form for an unfolding family $X\to S$ (e.g. [S1]) is a map defined by the period integrals of the primitive form, and is a multi-valent map from the complement of the discriminant divisor $S\setminus D$ of the base space $S$ of the family to its range domain (so called, period domain). In these classical cases [ibid], the range is isomorphic to the monodromy covering of $S\setminus D$, i.e. the quotient space of the universal covering $(S\setminus D)^{\sim}$ divided by the kernel of the monodromy representation of $\pi_{1}(S\setminus D,*)$. Then, the inversion problem for the period map (Siegel [Si] 1.13.) is to ask for a description of the inverse map from the range to the defining domain $S\setminus D$, and its answer were given by Eisenstein series (see [Si] Section13 and [S1] Section9)).
Recently, we studied what we call the virtual period integrals for some non-geometric primitive forms [S2], where the ranges are no-longer isomorphic to the monodromy coverings but are their quotients by additional groups which we call the winding group. Due to the existence of the winding group, a naive answer to the inversion problem for the virtual period maps can not exist.
The purpose of the present note is to illustrate such winding quotients by an elementary model of virtual period maps of rank 1. We observe that the winding group, acting on the monodromy covering space, acts also on the defining domain of the period map so that we obtain the winded quotient defining domain. Then, the classical inversion problem is reformulated in this new setting to a problem to describe the morphism from the winding quotient range to the winding quotient domain. We solve the new problem by a use of a $q$-multiplicatively periodic function $\mathcal{W}$ on the range whose pull-back to the monodromy covering (i.e. the winding covering of the range) is the Fourier expansion part of $\wp$-function up to a correction by Eisenstein series $E_{2}$ studied by Borcherds [B].
Remark 1.1.
Actually, the pull-back of the function $\mathcal{W}$ to the winding covering space has been studied in quite different contexts in mathematical physics either as the propagator $P_{0}^{\infty}(w_{1},w_{2};\tau,\overline{\tau})$ of BCOV theory by Si Li ([L1] Lemma 3.1, [L2] section 4 and [L-Z] 3.3) and also the propagator $P(z)$ by Dijkgraaf [D] (4.44) on elliptic curves, or as the Bergman kernel function $B(z_{1},z_{2})$ for the case genus =1 in the topological recursion by Eynard [E] p.339 Examples (c.f. also Tyurin [T] p.185 Example. $\omega_{\alpha}$ and Fay [F] (46)). We do not know yet whether such connection of the winding phenomenon with mathmatical physics is common for non-geometric virtual period maps in general or it remains only in these test examples. We need to test further cases [S2, G-S].
Notation. (1) We denote the set $\mathbb{C}\setminus\{0\}$ by $\mathbb{C}^{\times}$ because of its multiplicative structure.
(2) When the complex plane $\mathbb{C}$, the complex upper half plane $\mathbb{H}$ or the unit disc $\mathbb{D}$ is equipped with a particular coordinate, say $z$, then, in order to distinguish it from the one equipped with other coordinates, we put to $\mathbb{C}$, $\mathbb{H}$ or $\mathbb{D}$ the coordinate name as a subscript, e.g. $\mathbb{C}_{z}$, $\mathbb{H}_{z}$ or $\mathbb{D}_{z}$.
2. The virtual period map $z^{\tau}$ on $\mathbb{C}_{z}^{\times}$
We start to define a virtual period map of rank 1 by the solution of period equation.  [^1] We used in [S2] the term “virtual” for the period integrals whose associated root lattices do not have algebraic geometry origin. We employed in the present note the same term “virtual”. However, we use here the term “virtual” simply in the sense that the exponent $\tau$ of the period equation are non-real complex numbers.
Consider an ordinary differential equation in one variable $z$ and the unknown variable $w$:
[…]
where $\tau$ is a complex number such that $\rm{Im}(\tau)>0$. We call the equation the virtual period equation of rank 1. The solutions of the equation are, up to a constant factor, equal to the function $w=z^{\tau}=\exp\!\big(h\log(z)\big)$. We say that the solution $z^{\tau}$ “defines”, what we call, the virtual period map:
$z^{\tau}\ :\ \mathbb{C}_{z}^{\times}\ \longrightarrow\ \mathbb{C}_{w}^{\times},\ \ z\ \mapsto\ w=z^{\tau}$
from the punctured $z$-plane $\mathbb{C}_{z}^{\times}$ as the domain of the map to the punctured plane $\mathbb{C}_{w}^{\times}$ as the range of the map. What does this mean exactly geometrically? Obviously it is not a univalent map. So we define the map geometrically as follows. We first choose $1\in\mathbb{C}_{z}^{\times}$ as the base point in the domain, and define the value of the map $z^{\tau}$ at $1\in\mathbb{C}_{z}^{\times}$ to be 1. Then, the values of the map are defined by analytic continuations of $\exp\!\big(h\log(z)\big)$ along paths on the $z$-plane $\mathbb{C}_{z}^{\times}$ starting at 1. The monodromy of this map by the counter clock-wise turn around 0 of the $z$-variable is given by the multiplication of $\exp{\big(2\pi\sqrt{-1}\tau\big)}$ on $z^{\tau}$, i.e. its multiplication in the $w$-plane $\mathbb{C}_{w}^{\times}$. Since $|\exp{\big(2\pi\sqrt{-1}\tau\big)}|\!<\!1$, the monodromy group $\big(\exp{(2\pi\sqrt{-1}\tau)}\big)^{\mathbb{Z}}$ is an infinite cyclic group and gives a faithful representation of the fundamental group $\pi_{1}(\mathbb{C}_{z}^{\times},1)\simeq\mathbb{Z}$. Then the monodromy covering space of the defining domain of the map $z^{\tau}$ is the universal covering space $\mathbb{C}_{\mathrm{Log}z}$ of $\mathbb{C}_{z}^{\times}$, where we denote by $\mathrm{Log}z$ the coordinate of the universal covering space such that $\exp(\mathrm{Log}z)=z$. On the space $\mathbb{C}_{\rm{Log}z}$, we have the covering transformation group action
$\Gamma_{g}\ :=\ 2\pi\sqrt{-1}\mathbb{Z}:\qquad\mathrm{Log}z\mapsto\mathrm{Log}z+2\pi\sqrt{-1}n\quad\ \ for\ \ n\in\mathbb{Z}\qquad\qquad$
whose quotient map is the covering map $\exp:\mathbb{C}_{\mathrm{Log}z}\to\mathbb{C}_{z}^{\times}$.
3. Lifting of the period map $z^{\tau}$ to the monodromy covering space
Recall that $\mathrm{Log}z$ is the coordinate of the monodromy covering plane of $\mathbb{C}_{z}^{\times}$. Then, the virtual period map $z^{\tau}$ is lifted as the univalent map $w=\exp\big(\tau\mathrm{Log}z\big)$ on the $\mathrm{Log}z$-plane $\mathbb{C}_{\mathrm{Log}z}$. So the range of virtual period map $z^{\tau}$ means the range of this lifted map. Thus, we obtain the diagram
$\begin{array}[]{cccccc}\vskip 5.69046pt\qquad\mathrm{Log}z\ \in\!&\mathbb{C}_{\mathrm{Log}z}\!&\overset{\exp(\tau\mathrm{Log}z)}{-\!\!\!-\!\!\!-\!\!\!-\!\!\!\longrightarrow}&\mathbb{C}_{w}^{\times}\ \ni\ w=z^{\tau}=\exp(\tau{\mathrm{L}og}z)\\ \vskip 5.69046pt\exp(\mathrm{Log}z)\!\!\!\!\!\!\!\!\!\!&\downarrow&\quad\nearrow_{z^{\tau}}\!\!\!\!\!\!\!\!\!\!\\ \qquad z\ \in\!\!\!\!\!\!\!\!\!&\!\mathbb{C}_{z}\setminus\{0\}\!\!\!\!\!\!\!\!\!\!\end{array}$
We warn that the anti-diagonal arrow for the virtual period map $z^{\tau}$ is a multi-valent map so that the diagram is not commutative in the usual sense.
Non-existence of the inversion map? The naive “invertion problem” in this setting is to give an inversion map from $\mathbb{C}_{w}^{\times}$ to $\mathbb{C}_{z}^{\times}$ of the map $z^{\tau}$. However, the observation is that there does not exists such univalent holomorphic map which makes the diagram commutative, since the analytic continuation of any local inversion map becomes automatically the multi-valued function $z=w^{1/\tau}$ on $\mathbb{C}_{w}^{\times}$. That is, a naive solution of the classical inversion problem for the virtual period map $z^{\tau}$ by a univalent map from the range to the domain does not exist.
4. Winding of the monodromy covering space $\mathbb{C}_{\mathrm{Log}z}$ to the range $\mathbb{C}_{w}^{\times}$ of $z^{\tau}$
We saw that the virtual period map $z^{\tau}$ is presented by the univalent function $w=\exp\big(\tau\mathrm{Log}z\big)$ defined on the monodromy covering space $\mathbb{C}_{\mathrm{Log}z}$. Then, we immediately observe that its range $\mathbb{C}_{w}$ is the quotient of the monodromy covering space $\mathbb{C}_{\mathrm{Log}z}$ by the translation group action
$\Gamma_{w}\ :=\ \frac{2\pi\sqrt{-1}}{\tau}\mathbb{Z}:\qquad\ \mathrm{Log}z\mapsto\mathrm{Log}z+\frac{2\pi\sqrt{-1}}{\tau}n\quad\ for\ \ n\in\mathbb{Z}.\qquad\qquad$
Let us call the group $\Gamma_{w}$ the winding group and the map $\mathbb{C}_{\mathrm{Log}z}\to\mathbb{C}_{w}^{\times}$ the winding quotient map. In the other words, the range $\mathbb{C}_{w}^{\times}$ of $z^{\tau}$ is the winding quotient space of the monodromy covering space $\mathbb{C}_{\mathrm{Log}z}$ by the winding group $\Gamma_{w}$-action.
The key observation is that the group $\Gamma_{w}$ is caused by the group of the “symmetry” of the virtual period map $z^{\tau}=\exp(\tau\mathrm{Log}z)$. Namely, the winding group action leaves the function $z^{\tau}$ invariant:
$\Big(e^{\frac{2\pi\sqrt{-1}}{\tau}n}z\Big)^{\tau}=z^{\tau}\quad\text{for }n\in\mathbb{Z}$
or equivalently
$\exp\Big(\tau(\mathrm{Log}z+\frac{2\pi\sqrt{-1}}{\tau}n)\Big)=\exp\Big(\tau\mathrm{Log}z\Big)\quad\text{for }n\in\mathbb{Z}.$
In the other words, $\Gamma_{w}$ does not change the branch of $z^{\tau}$, whereas $\Gamma_{g}$ does change it.
5. Winding of the domain $\mathbb{C}_{z}^{\times}$ of $z^{\tau}$ to the elliptic curve $E_{\tau}$
In previous sections, we observed that the monodromy covering space $\mathbb{C}_{\mathrm{Log}z}$ obtained two different group actions, one: winding group $\Gamma_{w}$ and the other: the monodromy covering group $\Gamma_{g}$. We observe immediately that the two actions commute to each other. Consequently, the winding group $\Gamma_{w}$ acts also on the defining domain $\mathbb{C}_{z}^{\times}$ of the map $z^{\tau}$ and the covering group $\Gamma_{g}$ acts on the range $\mathbb{C}_{w}^{\times}$. Thus we obtain the following commutative Cartesian diagram:
[…]
where $E_{\tau}$ is the quotient of the domain $\mathbb{C}_{z}^{\times}$ by the winding group $\Gamma_{w}$-action. Actually, $E_{\tau}$ is an elliptic curve given by the quotient of the plane $\mathbb{C}_{\mathrm{Log}z}$ by the action of the lattice given by
$\Gamma_{g}+\Gamma_{w}\ =\ \frac{2\pi\sqrt{-1}}{\tau}(\mathbb{Z}+\mathbb{Z}\tau).$
The virtual period map $z^{\tau}$ induces the new map $E_{\tau}\to\mathbb{C}_{w}^{\times}$ (recall that $z^{\tau}$ is invariant under the action of $\Gamma_{w}$) so that $\mathbb{C}_{w}^{\times}$ becomes the monodromy covering space of $E_{\tau}$ for the new map, since $\pi_{1}(E_{\tau},0)\simeq\Gamma_{g}+\Gamma_{w}$ and the kernel of the new monodromy presentation is equal to $\Gamma_{w}$.
In this new situation for the virtual period map, what does the “inversion problem” may mean? One answer, which we shall employ in the present note, is that it is a question asking to describe the inverse morphism $\mathbb{C}_{w}^{\times}\to E_{\tau}$, i.e. the vertical down arrow in the RHS of the diagram ()(*). Namely, for a prescribed $w\in\mathbb{C}_{w}^{\times}$, we want to find the unique point, say $e$, in $E_{\tau}$ whose inverse image in $\mathbb{C}_{w}^{\times}$ (=the set of virtual periods for $e$) contains $w$, and to describe $e$ in terms of $w$.
In order to realize this program, we need to fix a coordinate presentation of the point $e\in E_{\tau}$. It is well-known that the elliptic curve $E_{\tau}$ is realized as the plane curve:
$\qquad\qquad E_{\tau}\ \subset\ \mathbb{P}^{2},\quad\ u\bmod\mathbb{Z}+\mathbb{Z}\tau\ \ \mapsto\ \ (\wp(u,\tau):\wp^{\prime}(u,\tau):1)$
by the use of Weierstrass $\wp$-function $\wp(u,\tau)$ and its derivative $\wp^{\prime}(u,\tau)$ given in the coordinate
$u=\frac{\tau}{2\pi\sqrt{-1}}\mathrm{Log}z$
where the scaling factor $\frac{\tau}{2\pi\sqrt{-1}}$ comes from the scaling of the group $\Gamma_{g}+\Gamma_{w}=\frac{2\pi\sqrt{-1}}{\tau}(\mathbb{Z}+\mathbb{Z}\tau)$ relative to the lattice $\mathbb{Z}+\mathbb{Z}\tau$ for the $\wp$-function. Thus, the new inversion problem asks to describe the pair coordinates $(\wp(u,\tau),\wp^{\prime}(u,\tau))$ of the point $e\in E_{\tau}$ in terms of $w=\exp(\tau\mathrm{Log}z)$ and $\tau$.  [^2] The realization of $E_{\tau}$ as a plane curve is not unique, but there are severals (e.g. [S2]), and in the present note, we have chosen just one of them, i.e. $\mathrm{A}_{2}$-type realization. It would be also interesting to study the other cases.
6. $q$-multiplicatively periodic function $\mathcal{W}(w,q)$
In the present section, we study a meromorphic function $\mathcal{W}$ on $\mathbb{C}_{w}^{\times}$ which is invariant by the multiplicative action of $q^{\mathbb{Z}}:=\exp(2\pi\sqrt{-1}\tau\mathbb{Z})$ in order to answer to the new inversion problem.
Following the classical idea of the construction of the $\wp$-function by its partial fractional expantion on the $u$-plane $\mathbb{C}_{u}$ (which is additively doubly periodic, having second order poles at $\mathbb{Z}+\mathbb{Z}\tau$), we construct a meromorphic function $\mathcal{W}$ on $w\in\mathbb{C}_{w}^{\times}$ which is multiplicatively periodic in $w$ by the periods $q^{\mathbb{Z}}$, i.e. $\mathcal{W}(wq,q)=\mathcal{W}(w,q)$, having second order poles on the set $\cup_{m\in\mathbb{Z}}\{1-wq^{m}=0\}$.
Let us consider an infinite sum of rational functions on $\mathbb{C}_{w}^{\times}\times\mathbb{D}_{q}^{\times}$:
$(1)\qquad\qquad\qquad\qquad\qquad\qquad\qquad\mathcal{W}(w,q)\ \ :=\ \ \sum_{m\in\mathbb{Z}}\frac{wq^{m}}{(1-wq^{m})^{2}}\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad$
whose $m$-th term $\frac{wq^{m}}{(1-wq^{m})^{2}}$ is a rational function having the second order pole at $w=q^{-m}$ which is invariant by the involution: $(w,q)\mapsto(w^{-1},q^{-1})$.
Assertion A. The sum obtained by removing the $m$-th term from $(1)$ converges uniformly on any compact subset of the domain $\mathcal{U}_{m}:=\{(w,q)\in\mathbb{C}_{w}^{\times}\times\mathbb{D}_{q}^{\times}\mid|q|^{-m+1}<|w|<|q|^{-m-1}\}$. So the sum $(1)$ defines a meromorphic function on $\mathcal{U}_{m}$. Since we have the open covering $\mathbb{C}_{w}^{\times}\times\mathbb{D}_{q}^{\times}=\cup_{m\in\mathbb{Z}}\ \mathcal{U}_{m}$, the sum $(1)$ converges to a meromorphic function on $\mathbb{C}_{w}^{\times}\times\mathbb{D}_{q}^{\times}$, which we shall denote again by $\mathcal{W}(w,q)$. The function $\mathcal{W}(w,q)$ is $q$-multiplicatively periodic.
$(2)\qquad\qquad\qquad\qquad\qquad\qquad\qquad\mathcal{W}(wq,q)\ \ =\ \ \mathcal{W}(w,q)\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad$
Proof.
We first decompose the sum in two parts:
$(3)\qquad\qquad\qquad\mathcal{W}(w,q)-\frac{wq^{m}}{(1-wq^{m})^{2}}\ =\ \ \sum_{n\in\mathbb{Z}_{>m}}\frac{wq^{n}}{(1-wq^{n})^{2}}+\sum_{n\in\mathbb{Z}_{>-m}}\frac{w^{-1}q^{n}}{(1-w^{-1}q^{n})^{2}}\qquad\qquad\qquad\qquad\qquad$
where, using the involutivity, the second part is: $\sum_{n\in\mathbb{Z}_{<m}}\frac{wq^{n}}{(1-wq^{n})^{2}}=\sum_{n\in\mathbb{Z}_{>-m}}\frac{w^{-1}q^{n}}{(1-w^{-1}q^{n})^{2}}$.
Notation. For any compact subset $K$ of the domain $\mathcal{U}_{m}$ and for any complex valued continuous function $f$ on $\mathcal{U}_{m}$, we set $f|\!|_{K}:=\mathrm{sup}\{|f(w,q)|\mid(w,q)\in K\}$.
We next show, for the given $m\in\mathbb{Z}$ and $K$, there exists a positive constants $C\in\mathbb{R}_{>0}$ such that
[…]
Then, we set
$C:=\mathrm{sup}\{2,\ |\!|\frac{1}{1-wq^{n}}|\!|_{K}\ \text{for}\ m<n<N,\ |\!|\frac{1}{1-w^{-1}q^{n}}|\!|_{K}\ \text{for}\ -m<n<N\ \}\qquad\Box\ )$
Thus, we finally have the following majorations for each parts of $(3)$
[…]
which implies the compact uniform convergence of the sum in the right hand side of $(3)$ so that the sum gives a holomorphic function on the domain $\mathcal{U}_{m}$.
The function $\mathcal{W}(wq,q)$ is defined by one shift of the summation index of the sum $(1)$. The compact uniform convergence of the sum (1) implies that we can change the summation order, implying the equality (2) as meromorphic functions. ∎
Assertion B. The function $\mathcal{W}$ and the $\wp$-function are related by the following pull-back relations.
$\begin{array}[]{ccl}\wp(u,\tau)&=&(2\pi\sqrt{-1})^{2}\ \mathcal{W}(w,q)\ -\ \frac{\pi^{2}}{3}E_{2}(\tau)\\ \wp^{\prime}(u,\tau)&=&(2\pi\sqrt{-1})^{3}\ \mathcal{W}^{\prime}(w,q)\end{array}$
where we set $w=\exp(2\pi\sqrt{-1}u),\ q=\exp(2\pi\sqrt{-1}\tau)$, and $E_{2}(\tau)=\frac{1}{\pi^{2}}\sum_{m}\big(\sum_{n}^{\prime}\frac{1}{(m\tau+n)^{2}}\big)$ is the Eisenstein series of weight 2 (here $\sum_{n}^{\prime}$ means n=0n\!=\!0 is omitted if m=0m\!=\!0, see Zagier [Z] p. 19).
Proof.
For each fixed $\tau\in\mathbb{H}$, the substitution $\mathcal{W}(\exp(2\pi\sqrt{-1}u),\exp(2\pi\sqrt{-1}\tau))$ is a doubly periodic meromorphic function in $u$ having poles at the lattice $\mathbb{Z}+\mathbb{Z}\tau$, and the principal part of its Laurent expansion at $u=0$ is equal to $\frac{1}{(2\pi\sqrt{-1})^{2}}\frac{1}{u^{2}}$. Then, the difference of the $\wp(u,\tau)$ and the substitution $(2\pi\sqrt{-1})^{2}\mathcal{W}(\exp(2\pi\sqrt{-1}u),\exp(2\pi\sqrt{-1}\tau))$ is a doubly periodic function on $u\in\mathbb{C}_{u}$ without poles so that it is a constant function on $u$. That is, the difference is a function depending only on $\tau$.
Actually, the differnce was already calculated by R. Borcherds as follows. In [B] Chapter7, Borcherds give the following Fourier expansion of the $\wp$-function (see also Aoki [A] (3.10)).
$\wp(u,\tau)=(2\pi\sqrt{-1})^{2}\sum_{m\in\mathbb{Z}}\sum_{n\in\mathbb{Z}_{>0}}n\exp\big(\pm 2\pi\sqrt{-1}n(u+m\tau)\big)-\frac{\pi^{2}}{3}E_{2}(\tau)$
whenever $\Im(u+m\tau)$ is non-zero for all integers $m$, where the sign $\pm$ means ++ if $|\exp(2\pi\sqrt{-1}n(u+m\tau))|>1$ and - if $|\exp(2\pi\sqrt{-1}n(u+m\tau))|<1$. That is, the formula holds on the open set $\mathbb{C}_{u}\times\mathbb{H}_{\tau}\setminus\cup_{m\in\mathbb{Z}}\{\Im(u+m\tau)=0\}$. Clearly, it projects onto the defining domain $\mathbb{H}_{\tau}$ of $E_{2}(\tau)$. ∎
Remark 6.1.
The $q$-multiplicative periodicity of $\mathcal{W}(w,q)$ comes from the monodromy group $\Gamma_{g}$ action on the covering space $\mathbb{C}_{\rm{Log}z}$ (whose invariant is the exponential function) so that one should not confuse it with the winding group $\Gamma_{w}$ invariance. Symbolically speaking, $\mathcal{W}(w,q)$ is the winding quotient of the exponential map $\exp:\mathbb{C}_{Loz}\to\mathbb{C}_{z}^{\times}$, i.e. $\mathcal{W}=\exp\bmod\ \Gamma_{w}$. In the other words, the function $\mathcal{W}$ is the shadow of the exponential function by the winding group $\Gamma_{w}$.
Remark 6.2.
Yosuke Ohyama has informed author that the function $\mathcal{W}(w,q)$ can be expressed by a bilateral basic hypergeometric series ([G-R] chap.5.) as follows;
$_{2}\psi_{2}\begin{bmatrix}\vskip-8.5359ptw,w&\\ \vskip-5.69046pt&\!\!;q,q\\ wq,wq\!\!&\end{bmatrix}=\sum_{n=-\infty}^{\infty}\frac{(w,w;q)_{n}}{(wq,wq;q)_{n}}q^{n}=\sum_{n=-\infty}^{\infty}\frac{(1-w)^{2}}{(1-wq^{n})^{2}}q^{n}=\frac{(1-w)^{2}}{w}\mathcal{W}(w,q).$
7. Solution to the new inversion problem
The descriptions of the morphism: $\mathbb{C}_{w}^{\times}\to E_{\tau}\subset\mathbb{P}^{2}$ is achieved as follows. This gives an answer to the new inversion problem posed at the end of section 5.
Theorem 7.1.
The pair of meromorphic functions in $w\in\mathbb{C}_{w}^{\times}$
$\begin{array}[]{ccl}(\ 2\pi\sqrt{-1})^{2}\mathcal{W}(w,\exp(2\pi\sqrt{-1}\tau)-\frac{\pi^{2}}{3}E_{2}(\tau)\ :\ (2\pi\sqrt{-1})^{3}\mathcal{W}^{\prime}(w,\exp(2\pi\sqrt{-1}\tau)\ :\ 1\ )\end{array}$
gives a morphism $\mathbb{C}_{w}^{\times}\to E_{\tau}\subset\mathbb{P}^{2}$, making the diagram ()(*) in section 5 commutative.
Proof.
This follows from Assertion B. in section 6. ∎
Acknowledgement: The author express his gratitude to Hiroki Aoki who informed him the Borcherds work and helped him in solving the new inversion problem, and to Akishi Ikeda who explained him the roles of the pull-back of the function $\mathcal{W}$ to the winding covering space in mathematical physics and informed him its related literatures (see Remark 1.1). He also express his gratitude to Masahiko Yoshinaga for discussions during the preparation of the present note.